OpenStudy (anonymous):

No idea how to solve? Solve 4^x = 12

jimthompson5910 (jim_thompson5910):

Hint: \[\Large b^x = y\] converts to \[\Large \log_{b}(y) = x\]

OpenStudy (anonymous):

okay...

jimthompson5910 (jim_thompson5910):

does that make sense

OpenStudy (anonymous):

or use \(b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\) if you want a decimal approximation

OpenStudy (anonymous):

Kinda, but could you maybe show me the steps you would take to get this answer. Sorry I grasp it better that way :/

jimthompson5910 (jim_thompson5910):

Similar example: Solve 5^x = 13 for x \[\Large 5^x = 13\] \[\Large \log_{5}(13) = x\] \[\Large x = \log_{5}(13)\] \[\Large x = \frac{\ln(13)}{\ln(5)}\] \[\Large x \approx 1.59369264\]

jimthompson5910 (jim_thompson5910):

That's not the answer to your problem, but it shows you how do these type of problems.

OpenStudy (anonymous):

Got it thanks! But what if its like: 3^x – 4 = 7^x + 9

jimthompson5910 (jim_thompson5910):

are x-4 and x+9 exponents?

OpenStudy (anonymous):

Yes!

jimthompson5910 (jim_thompson5910):

is the problem \[\Large 3^{x-4} = 7^{x+9}\]

OpenStudy (anonymous):

Yes :)

jimthompson5910 (jim_thompson5910):

alright, one moment

jimthompson5910 (jim_thompson5910):

Similar Example: Solve \(\Large 7^{x-2} = 5^{x+6}\) for x \[\Large 7^{x-2} = 5^{x+6}\] \[\Large \ln\left(7^{x-2}\right) = \ln\left(5^{x+6}\right)\] \[\Large (x-2)\ln(7) = (x+6)\ln(5)\] \[\Large x*\ln(7)-2\ln(7) = x*\ln(5)+6\ln(5)\] \[\Large x*\ln(7) = x*\ln(5)+6\ln(5)+2\ln(7)\] \[\Large x*\ln(7)-x*\ln(5) = 6\ln(5)+2\ln(7)\] \[\Large x\left(\ln(7)-\ln(5)\right) = 6\ln(5)+2\ln(7)\] \[\Large x = \frac{6\ln(5)+2\ln(7)}{\ln(7)-\ln(5)}\] \[\Large x \approx 40.2661684921349\] Note: the last line is the approximate solution and the second to last line is the exact solution This isn't the answer to your problem, but it shows you how to solve problems like this.

OpenStudy (anonymous):

Thanks!

jimthompson5910 (jim_thompson5910):

you're welcome